SpaceRef

SpaceRef


NASA Hubble Space Telescope Daily Report #4141

Status Report From: Space Telescope Science Institute
Posted: Friday, June 23, 2006

image

HUBBLE SPACE TELESCOPE - Continuing to collect World Class Science

DAILY REPORT # 4141

PERIOD COVERED: UT June 22, 2006 (DOY 173)

OBSERVATIONS SCHEDULED

NIC3 10632

Searching for galaxies at z>6.5 in the Hubble Ultra Deep Field

We propose to obtain deep ACS {F606W, F775W, F850LP} imaging in the area of the original Hubble Ultra Deep Field NICMOS parallel fields and - through simultaneous parallel observations - deep NICMOS {F110W, F160W} imaging of the ACS UDF area. Matching the extreme imaging depth in the optical and near-IR bands will result in seven fields with sufficiently sensitive multiband data to detect the expected typical galaxies at z=7 and 8. Presently no such a field exist. Our combined optical and near-IR ultradeep fields will be in three areas separated by about 20 comoving Mpc at z=7. This will allow us to give a first assessment of the degree of cosmic variance. If reionization is a process extending over a large redshift interval and the luminosity function doesn't evolve strongly beyond z=6, these data will allow us to identify of the order of a dozen galaxies at 6.5 < z < 8.5 - using the Lyman break technique - and to place a first constrain on the luminosity function at z>6.5. Conversely, finding fewer objects would be an indication that the bulk of reionization is done by galaxies at z=6. By spending 204 orbits of prime HST time we will capitalize on the investment of 544 prime orbits already made on the Hubble Ultra Deep Field {UDF}. We have verified that the program as proposed is schedulable and that it will remain so even if forced to execute in the 2-gyro mode. The data will be non-proprietary and the reduced images will be made public within 2 months from the completion of the observations.

NIC1/NIC2/NIC3 8793

NICMOS Post-SAA calibration - CR Persistence Part 4

A new procedure proposed to alleviate the CR-persistence problem of NICMOS. Dark frames will be obtained immediately upon exiting the SAA contour 23, and every time a NICMOS exposure is scheduled within 50 minutes of coming out of the SAA. The darks will be obtained in parallel in all three NICMOS Cameras. The POST-SAA darks will be non-standard reference files available to users with a USEAFTER date/time mark. The keyword 'USEAFTER=date/time' will also be added to the header of each POST-SAA DARK frame. The keyword must be populated with the time, in addition to the date, because HST crosses the SAA ~8 times per day so each POST-SAA DARK will need to have the appropriate time specified, for users to identify the ones they need. Both the raw and processed images will be archived as POST-SAA DARKSs. Generally we expect that all NICMOS science/calibration observations started within 50 minutes of leaving an SAA will need such maps to remove the CR persistence from the science images. Each observation will need its own CRMAP, as different SAA passages leave different imprints on the NICMOS detectors.

NICS3 10754

Spectrophotometric Standards for Cross-Observatory Calibration

This program will obtain NICMOS spectrophotometry of four main sequence A stars and four K giants, each selected from the Spitzer IRAC photometric calibration target and/or candidate calibration target lists. These observations will supplement existing HST observations of DA white dwarfs and solar analogs, and will provide a broad base of stellar types for spectrophotometric cross calibration of HST, Spitzer, and eventually JWST. The targets are chosen to be faint enough for unsaturated observations with JWST NIRSPEC, yet still bright enough for high signal to noise in relatively short observations with HST+NICMOS and with Spitzer+IRAC.

FLIGHT OPERATIONS SUMMARY:

Significant Spacecraft Anomalies: (The following are preliminary reports of potential non-nominal performance that will be investigated.)

HSTARS: (None)

COMPLETED OPS REQUEST:

17802-0 - ACS Recovery to Operate1 @ 173/21:11z

COMPLETED OPS NOTES: (None)

                                  SCHEDULED     SUCCESSFUL 
FGS GSacq                      10                      10 
FGS REacq                       02                      02 
OBAD with Maneuver          26                      26 

SIGNIFICANT EVENTS:

ACS Transition to Operate1

Ops Request 17802-0 was completed at 173/21:11:12, successfully transitioning ACS from Suspend to its Operate1 state. In this state, ACS normal engineering data collection can be observed.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

SpaceRef Newsletter